### v2

# On the closure of a plane ray that limits onto itself

Abstract.
We show that the closure of any self-entwined ray in the plane must contain a Cantor set of mutually disjoint continua. This is false in dimension three.
Key words and phrases:
self-entwined, non-Suslinian, ray, indecomposable, plane topology
2020 Mathematics Subject Classification:
54F15, 37B45
1. Introduction
In this paper we study a planar form of entanglement or recurrence which gives rise to a complicated structure within the surrounding space. Specifically we examine properties of planar continua which contain self-entwined rays (i.e. rays that limit onto themselves). We show that these continua have uncountable structures.
1.1. Terminology
A ray is a continuous one-to-one image of $0,\infty)$. A ray is self-entwined if no open set intersects it in an arc. This means that each point is revisited to within $\varepsilon$ by points further along the ray (see Section 2.2). A continuum is a compact connected metric space. A continuum is indecomposable if it cannot be written as a union of two proper subcontinua. A continuum is non-Suslinian if it contains an uncountable collection of pairwise-disjoint, non-degenerate subcontinua.
Indecomposable continua are known to be non-Suslinian (cf. [8, §1.1]).
1.2. Motivation and main result
Self-entwined rays are commonly found in continua which arise from folding algorithms and chaotic systems. For example, they exist within the arc components of Plykin and Ikeda attractors (see Figures 1 and 2). These planar attractors are non-Suslinian because they are indecomposable; they even contain neighborhoods that are homeomorphic to the Cantor fence $2^{\omega}\times[0,1]$.
The goal of this paper is to establish the general result:
Theorem 1.
If $X$ is a self-entwined ray in $\mathbb{R}^{2}$, then $\overline{X}$ is non-Suslinian.
Although every ray is a rational space, meaning that it has a basis of open sets with countable boundaries, Theorem 1 implies that the closure of a self-entwined plane ray cannot be rational (cf. [8, §1.3]).
1.3. Context and limitations
Jones [4, 5] proved that a plane ray cannot be locally connected unless it is locally compact. From [7] it follows that each locally connected plane ray is homeomorphic to either $0,\infty)$, the unit circle $S^{1}$, or $S^{1}\cup[0,1]$. In particular, planar self-entwined rays are not locally connected.
Tymchatyn [12, Example 3] constructed a rational continuum which is the closure of a self-entwined ray. So Theorem 1 is false outside the plane. Tymchatyn’s continuum, which can be embedded into $\mathbb{R}^{3}$, was already known to be non-planar because it is hereditarily locally connected but not finitely-Suslinian [8]. Moreover, the ray in this continuum is locally connected.
Curry [2] showed that the closure of a self-entwined plane ray is an indecomposable continuum provided that it is $1$-dimensional and non-separating or finitely separating. On the other hand, the Sierpiński carpet, which is $1$-dimensional and locally connected with infinitely-many complementary components, is easily seen to contain a dense self-entwined ray. So the non-Suslinian property cannot always be derived from Curry’s result.
Figure 1. The Ikeda map f(z)=1+0.9zexp(i(0.4-61+|z|2))        produces a strange attractor $K$ which contains a fixed point $p\approx 1.114-2.285i$. The arc component of $p$ in $K$ is a self-entwined ray with initial point $p$. Iterates of I={x-2.285i:0≤x≤1.114}        approximate initial segments of the ray since $f$ is injective on II, and $I\setminus\{p\}$ is contained in the basin of attraction for $K$. The arcs $f^{6}[I]$ and $f^{7}[I]$ are shown above.
Figure 2. A planar Plykin attractor $P$ is generated by the spherical equations of [6], followed by a stereographic projection from the point $\langle 1/\sqrt{2},0,1/\sqrt{2}\rangle$. As with the Ikeda attractor, approximations of a self-entwined ray in $P$ are obtained by iteration on an arc $I$ which ends at a fixed point of the attractor (in planar coordinates the fixed point is $p\approx\langle 1.26,-1.53\rangle$). This self-entwined ray is remarkable in that it has bounded curvature.
The images in Figures 1 and 2 were created in Python via precise numerical calculations.
2. Preliminaries
2.1. Arcs of rays
An arc is a topological copy of the interval $[0,1]$. If $X$ is any ray with an associated mapping $f$, and $a,b]\subset[0,\infty)$ is a closed and bounded interval, then $f\restriction[a,b]$ is a homeomorphism and $f[a,b]$ is an arc in $X$. By an arc of the ray $X$, we mean an arc of the form $f[a,b]$. It is a consequence of Sierpiński’s theorem [11, Theorem 5.16] that if $X$ is not compact then every continuum in $X$ is of this form; see [1, Lemma 2.2].
2.2. Definition of self-entwined
Below is an equivalent formulation of the self-entwined property in terms of the ray function.
Proposition 2.
If $X$ is a ray and $f$ is a continuous one-to-one mapping of $0,\infty)$ onto $X$, then the following are equivalent.
  • (a)
    $X$ is self-entwined;
  • (b)
    $X$ has no compact neighborhoods;
  • (c)
    each arc of $X$ has empty interior in $X$;
  • (d)
    for every $x\in X$ there exists a sequence $r_{n}\in[0,\infty)$ such that $r_{n}\to\infty$ and $f(r_{n})\to x$.
Proof.
[…]
2.3. Connectedness im-kleinen
A space $X$ is said to be connected im-kleinen at a point $x\in X$ if every neighborhood of $x$ contains a connected neighborhood of $x$.
The following is due to Fitzpatrick and Lelek [3].
Proposition 3.
Let $X$ be a continuum. If $X$ is Suslinian, then $X$ is connected im-kleinen at each point of a dense subset.
Idea of proof.
If $U$ is a non-empty open subset of $X$ that has only countably many non-degenerate components, then some component of $U$ must have interior in $X$ by the Baire category theorem. So if $X$ is Suslinian, it is possible to recursively define a decreasing sequence of continua with interiors in $X$, which intersect down to a single point of $X$ that is contained in each interior. ∎
2.4. Discs and theta curves
A $\theta$-curve is a space made of three arcs $\alpha_{1},\alpha_{2},\alpha_{3}$ which have the same endpoints $a$ and $b$ and are otherwise disjoint.
It is a standard fact of plane topology that one of the three arcs comprising a $\theta$-curve is separated from $\infty$ by the union of the other two. Another fact relevant to the proposition below is that if $F$ is a closed subset of $\mathbb{R}^{2}$ which separates two points $p$ and $q$ (that is, $p$ and $q$ belong to different components of $\mathbb{R}^{2}\setminus F$), then some component of $F$ separates $p$ and $q$. This is a well-known corollary of Janiszewski’s theorem.
Proposition 4.
Let $\theta=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}$ be a theta curve of arcs with endpoints $a$ and $b$, such that $\alpha_{1}\cup\alpha_{3}$ separates $\alpha_{2}$ from $\infty$. Let $D$ be a closed disc missing $a$ and $b$. If $p\in\alpha_{1}\cap D^{\mathrm{o}}$ and $q\in\alpha_{3}\cap D^{\mathrm{o}}$, then there exists an arc $\beta\subset\alpha_{2}\cap D$ which separates $p$ from $q$ in $D$, and is such that $\beta\cap\partial D$ consists only of the two endpoints of $\beta$.
Sketch of proof.
Let $\alpha_{4}$ be an arc that lies in the unbounded component of $\mathbb{R}^{2}\setminus(\alpha_{1}\cup\alpha_{3})$, which has endpoints $a$ and $b$, and is such that $\alpha_{4}\cap D=\varnothing$. The simple closed curve $\alpha_{2}\cup\alpha_{4}$ separates $p$ and $q$. It follows that $\alpha_{2}\cap D$ separates $p$ from $q$ in $D$. Thus some component $\gamma$ of $\alpha_{2}\cap D^{\mathrm{o}}$ separates $p$ from $q$ in $D^{\mathrm{o}}$. Let $\beta=\overline{\gamma}$ (see Figure 3).∎
Figure 3. Illustration for Proposition 4.
2.5. Notation
For any set $A\subset\mathbb{R}^{2}$ we let $\overline{A}$, $\partial A$, and $A^{\mathrm{o}}$ denote the closure, boundary, and interior of $A$, respectively, relative to the topology of $\mathbb{R}^{2}$. We denote by $2^{n}$ the set of functions from $\{0,\ldots,n-1\}$ to $\{0,1\}$. Thus an element $\sigma\in 2^{n}$ can be viewed as a binary sequence of length $n$. That is, $\sigma=\langle\sigma(0),\ldots,\sigma(n-1)\rangle$ where $\sigma(i)\in\{0,1\}$ for each $i<n$.
$2^{<\omega}=\bigcup_{n=0}^{\infty}2^{n}$
denotes the set of all finite binary sequences. Finally, $2^{\omega}$ is the set of all infinite binary sequences index by $\omega=\{0,1,2,3,\ldots\}$. If $\sigma\in 2^{<\omega}$, then $\sigma^{\frown}0$ and $\sigma^{\frown}1$ represent the one-element extensions of $\sigma$ obtained by appending 0 and $1$, respectively. If $\tau\in 2^{\omega}$ and $n<\omega$, then $\tau\restriction n$ is the sequence made of the first $n$ terms of $\tau$.
3. Rays that limit to arcs on both sides
Let $X$ be a ray in $\mathbb{R}^{2}$. Let $f$ be a continuous one-to-one mapping of $0,\infty)$ onto $X$. For each integer $n\geq 1$ let $S_{n}$ be a simple closed curve containing the arc $f[0,n]$. Let $U_{n}$ and $V_{n}$ be the components of $\mathbb{R}^{2}\setminus S_{n}$. Given $t>0$ and an integer $n>t$, we say that $X$ limits to $f(t)$ from both sides of $f[0,n]$ if $f(t)\in\overline{U_{n}\cap X}\cap\overline{V_{n}\cap X}$.
A point $x=f(t)\in X$ is a two-sided limit point of $X$ if there exists an integer $n>t$ such that $X$ limits to $f(t)$ from both sides of $f[0,n]$.
Let $A,B,C\subset\mathbb{R}^{2}$. An arc $\alpha\subset\mathbb{R}^{2}$ is $ABC$-minimal if $\alpha\cap A$ and $\alpha\cap C$ are the endpoints of $\alpha$, and $\alpha\cap B\neq\varnothing$. It is easy to prove that if $A$, $B$, and $C$ are closed and pairwise-disjoint, and $\alpha$ is any arc meeting all three sets, then $\alpha$ contains an arc which is $ABC$, BACBAC, or $ACB$-minimal.
Lemma 5.
If $X$ is self-entwined and $\overline{X}$ is connected im-kleinen at a dense set of points, then the set of two-sided limit points of $X$ is dense in $X$.
Proof.
Suppose that $X$ is self-entwined and $\overline{X}$ is connected im-kleinen at a dense set of points. Let $U$ be any non-empty open subset of $\mathbb{R}^{2}$ meeting $X$. Let $D_{1},D_{2},D_{3}$ be pairwise-disjoint closed discs in $U$ whose interiors intersect $\overline{X}$. For each $i\in\{1,2,3\}$ let $K_{i}\subset X$ be a continuum which has interior in $X$ and is contained in the interior of $D_{i}$. There are 7 pairwise-disjoint arcs in $X$ which meet every $K_{i}$. Each one contains an arc that is $K_{1}K_{2}K_{3}$, or $K_{2}K_{1}K_{3}$, or $K_{2}K_{3}K_{1}$-minimal. At least 3 of the 7 minimal arcs, say $\alpha_{1},\alpha_{2},\alpha_{3}$, are of the same type. By re-indexing the $D^{\prime}s$ and $K^{\prime}s$ we may assume that the three arcs are $K_{1}K_{2}K_{3}$-minimal. Upon identifying $K_{1}$ and $K_{3}$ with points $a$ and $b$, we have that $\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}$ is a $\theta$-curve. Let us assume that $\alpha_{2}$ is separated from $\infty$ by $\alpha_{1}\cup\alpha_{3}$. Choose $p\in\alpha_{1}\cap K_{2}$ and $q\in\alpha_{3}\cap K_{2}$. By Proposition 4, there exists an arc $\beta\subset\alpha_{2}\cap D_{2}$ which separates $p$ from $q$ in $D_{2}$, and is such that $\beta\cap\partial D_{2}$ consists of only the two endpoints of $\beta$. We claim that $\beta$ contains a two-sided limit point of $X$.
Let $n$ be an integer such that $\beta\subset f[0,n]$. Since $\beta$ is nowhere dense in $X$, $\beta$ is contained in the union of $\overline{U_{n}\cap X}$ and $\overline{V_{n}\cap X}$. If each of these sets intersects $\beta$, then by connectedness of $\beta$ there exists $x\in\overline{U_{n}\cap X}\cap\overline{V_{n}\cap X}\cap\beta.$ Then $x$ is a two-sided limit point of $X$ in $U$. Otherwise, one of the two sets misses $\beta$ which we will show leads to a contradiction. The contradiction will be that there exists an arc $\gamma$ that misses $K_{2}$, and separates $p$ and $q$ in $D_{2}$.
Without loss of generality, assume that $\overline{V_{n}\cap X}\cap\beta=\varnothing$. Let $G$ be the component of $\beta$ in $D_{2}\cap\overline{V_{n}}$. Since $\beta\cap\partial D_{2}$ consists of only the two endpoints of $\beta$, we have that $G$ is a closed topological disc and $\beta\subset\partial G$. Further, $G\setminus\overline{V_{n}\cap X}$ is a neighborhood of $\beta$ in the space $G$. Thus there exists an arc $\gamma\subset G\setminus\overline{V_{n}\cap X}$ which has the same endpoints as $\beta$ and is otherwise contained in $G^{\mathrm{o}}$. Choosing $\gamma$ close to $\beta$ (closer in Hausdorff distance than $d(p,\beta)$ and $d(q,\beta)$) will ensure that it separates $D_{2}$ between $p$ and $q$. Moreover, $\gamma\cap\overline{X}\subset\partial D_{2}$ by the fact $(G^{\mathrm{o}}\setminus\overline{V_{n}\cap X})\cap\overline{X}\subset(V_{n}\setminus\overline{V_{n}\cap X})\cap\overline{X}=\varnothing.$ Therefore $\gamma$ misses $K_{2}$, which contradicts that $K_{2}$ is connected. ∎
Remark 1.
If $\overline{X}$ is indecomposable, the set of two-sided limit points of $X$ may be empty. This is the case when $X$ is the accessible composant (arc component of the endpoint) of the Knaster buckethandle continuum.
4. Proof of Theorem 1
Let $X$ be a self-entwined ray in $\mathbb{R}^{2}$. Let $f\mathrel{\mathop{\ordinarycolon}}[0,\infty)\to X$ be a continuous one-to-one surjection. We want to show that $\overline{X}$ is non-Suslinian. This follows immediately from Proposition 3 if $\overline{X}$ is not connected im-kleinen at a dense set of points. For the remainder of the proof we may therefore assume that: $\overline{X}$ is connected im-kleinen at a dense set of points.
There exists a proper subcontinuum $K$ of $\overline{X}$ with interior in $\overline{X}$, and a disc $D\subset\mathbb{R}^{2}$ such that $D$ lies in the unbounded component of $\mathbb{R}^{2}\setminus K$ and $D^{\mathrm{o}}\cap X\neq\varnothing$. We will show that $\overline{X}$ contains an uncountable collection of pairwise-disjoint subcontinua stretching from $K$ to $D$.
By working in the unbounded component of $\mathbb{R}^{2}\setminus K$, we may assume that $K$ is non-separating. Hence if $\alpha$ is any $K(X\setminus K)K$-minimal arc, then $K\cup\alpha$ has exactly two complementary components. We may also assume that the ray’s endpoint $f(0)$ belongs to $K$, because there exists $f(t)\in K$ and $f[t,\infty)$ is a self-entwined ray with the same closure as $X$. Under this assumption, for every $x\in X\setminus K$ there exists a (unique) $K\{x\}K$-minimal arc in $X$. This arc is found by tracing $x$ to $K$, backwards and forwards along the ray.
For each $x\in X\setminus K$ we may now define:
  • $\alpha(x)$ to be the unique $K\{x\}K$-minimal arc of $X$,
  • $U(x)$ to be the bounded component of $\mathbb{R}^{2}\setminus(K\cup\alpha(x))$, and
  • $V(x)$ to be the unbounded component of $\mathbb{R}^{2}\setminus(K\cup\alpha(x))$.
Remark 2.
Note that $\alpha(x)\subset\partial U(x)\subset\alpha(x)\cup K$ and likewise for $V(x)$. Additionally, $\{U(x),\alpha(x),V(x)\}$ partitions the $K(X\setminus K)K$-minimal arcs of $X$, with the possible exception that $\alpha(x)$ may share an endpoint in $K$ with an arc in $U(x)$ or $V(x)$.
Observation 6.
If $x$ is any two-sided limit point of $X$ in $X\setminus K$, then $x\in\overline{U(x)\cap X}\cap\overline{V(x)\cap X}.$
Proof.
Suppose that $x=f(t)$ is a two-sided limit point in $X\setminus K$. Let $n>t$ be such that $X$ limits to $x$ from both sides of $f[0,n]$. There exists a closed domain $G$ such that $x\in G^{\mathrm{o}}$,
$G$ $\cap K=\varnothing,$
$G$ $\cap S_{n}\subset f(0,n)\text{, and}$
$G$ $\cap\alpha(x)=G\cap f[0,n]=\beta,$
where $\beta$ is an arc that contains $x$ and cuts $G$ into exactly two pieces. This is easy to see if one uses the fact that $f[0,n]$ is ambiently homeomorphic to a straight arc in which $f[0,n]\cap\alpha(x)$ is a neighborhood of $x$.
Let $G_{0}$ and $G_{1}$ be the components of $G\setminus\beta$. Since $G\setminus\beta\subset\mathbb{R}^{2}\setminus(K\cup\alpha(x))$ and $U(x)$ and $V(x)$ each intersect $G$, without loss of generality we may assume $G_{0}\subset U(x)$ and $G_{1}\subset V(x)$. Likewise, since $G\setminus\beta\subset\mathbb{R}^{2}\setminus S_{n}$ and $U_{n}$ and $V_{n}$ each intersect $G$, there exists $i\in\{0,1\}$ such that $G_{i}\subset U_{n}$ and $G_{1-i}\subset V_{n}$. Then $G\cap U_{n}=G_{i}$ and $G\cap V_{n}=G_{1-i}$. Now let $W$ be any open set containing $x$. Since $X$ limits to $x$ from both sides of $f[0,n]$, $W\cap G_{i}\cap X$ and $W\cap G_{1-i}\cap X$ are each non-empty. Therefore $W\cap U(x)\cap X$ and $W\cap V(x)\cap X$ are each non-empty. This shows that $x\in\overline{U(x)\cap X}\cap\overline{V(x)\cap X}$. ∎
We will now define a Cantor tree of compact regions of $\mathbb{R}^{2}$,
$\big\{D_{\sigma}\mathrel{\mathop{\ordinarycolon}}\sigma\in 2^{<\omega}\big\},$
and two-sided limit points $x_{\sigma}\in X$ so that:
  1. (1)
    $\alpha(x)\subset D_{\sigma}$ for all $x\in D_{\sigma}\cap X$,
  2. (2)
    $x_{\sigma^{\frown}i}\in(D\cap D_{\sigma})^{\mathrm{o}}$ for each $i\in\{0,1\}$, and
  3. (3)
    $\overline{U(x_{\sigma^{\frown}1})}\setminus K\subset U(x_{\sigma^{\frown}0})$.
Each $D_{\sigma}$ will be bounded by $K$ and finitely-many arcs of $X$.
First step: To begin, let $a\in D^{\mathrm{o}}\cap X$ be a two-sided limit point of $X$ provided by Lemma 5. By Observation 6 and Lemma 5 there exist two-sided limit points $b\in D^{\mathrm{o}}\cap U(a)\cap X$, $c\in D^{\mathrm{o}}\cap U(b)\cap X$, and $d\in D^{\mathrm{o}}\cap U(c)\cap X$. Define $D_{\varnothing}=\overline{U(a)}\setminus U(d)$ and put $x_{0}=b$ and $x_{1}=c$.
It is easy to see that (1) and (2) hold for $\sigma=\varnothing$. As for (3), $\overline{U(x_{1})}\setminus K\subset U(x_{1})\cup\alpha(x_{1})\subset U(x_{0})$.
Inductive step: Now suppose that $\sigma\in 2^{n}$ and $D_{\sigma}$ and $x_{\sigma^{\frown}0}$ and $x_{\sigma^{\frown}1}$ satisfy (1) through (3). Define
$D_{\sigma^{\frown}0}$ $=D_{\sigma}\cap\overline{V(x_{\sigma^{\frown}0})}$
$D_{\sigma^{\frown}1}$ $=D_{\sigma}\cap\overline{U(x_{\sigma^{\frown}1})}.$
By (2), Observation 6 and Lemma 5, there exists a two-sided limit point $x_{\sigma^{\frown}0,0}\in(D\cap D_{\sigma})^{\mathrm{o}}\cap V(x_{\sigma^{\frown}0})\cap X$. And there exists a two-sided limit point $x_{\sigma^{\frown}0,1}\in(D\cap D_{\sigma})^{\mathrm{o}}\cap V(x_{\sigma^{\frown}0})\cap U(x_{\sigma^{\frown}0,0})\cap X.$ Points $x_{\sigma^{\frown}1,0}$ and $x_{\sigma^{\frown}1,1}$ in $D_{\sigma^{\frown}1}$ can be found in the same way, with $U(x_{\sigma^{\frown}1})$ in the place of $V(x_{\sigma^{\frown}0})$. See Figures 4 and 5.
(1) through (3) now hold for all $\sigma\in 2^{n+1}$. (1) is verified by combining the condition on $D_{\sigma\restriction n}$ with the fact that $\overline{V(x_{\sigma^{\frown}0})}\cap X$ and $\overline{U(x_{\sigma^{\frown}1})}\cap X$ are unions of arcs $\alpha(x)$ (see Remark 2). (2) is obvious, and (3) holds for the reason given in the base case.
Figure 4. Construction of $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
Figure 5. Two possible arrangements of $D_{\sigma}$ ($\sigma\in 2^{2}$) within $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
For all $\sigma\in 2^{<\omega}$ we have:
  • $\blacksquare$
    $K\cup(D_{\sigma}\cap X)$ is a connected set that meets both $K$ and $D$ by (1) and (2),
  • $\blacksquare$
    $D_{\sigma^{\frown}0},D_{\sigma^{\frown}1}\subset D_{\sigma}$ by definition, and
  • $\blacksquare$
    $D_{\sigma^{\frown}0}\cap D_{\sigma^{\frown}1}\setminus K=\varnothing$ by (3).
For every $\tau\in 2^{\omega}$ define
$K_{\tau}=\bigcap_{n=0}^{\infty}\overline{K\cup(D_{\tau\restriction n}\cap X)}.$
Then $K_{\tau}$ is the intersection of a decreasing sequence of continua which contain $K$ and intersect $D$. Therefore $K_{\tau}$ is a continuum that contains $K$ and intersects $D$ [11, p.6].
Let $W$ be an open set containing $K$ such that $\overline{W}\cap D=\varnothing$. For each $\tau\in 2^{\omega}$ let $M_{\tau}$ be the component of some point of $K_{\tau}\cap D$ in $K_{\tau}\setminus W$. Note that different $\tau^{\prime}s$ produce disjoint $M_{\tau}$’s. And by the boundary bumping principle each $M_{\tau}$ is a non-degenerate subcontinuum of $\overline{X}$ [11, Theorem 5.4]. We conclude that $\{M_{\tau}\mathrel{\mathop{\ordinarycolon}}\tau\in 2^{\omega}\}$ is an uncountable collection of pairwise-disjoint, non-degenerate subcontinua of $\overline{X}$. Therefore $\overline{X}$ is non-Suslinian.
This completes the proof of Theorem 1. $\blacksquare$